OpenStudy (anonymous):

Solve using the zero factor property (w-4)(w-4)=0

OpenStudy (anonymous):

what is the zero factor property?

OpenStudy (anonymous):

w = 4 only

hero (hero):

Zero Product Property ab = 0

hero (hero):

But what you want to do in this case is the following steps: (w-4)(w-4) = 0 (w-4)^2 = 0 w - 4 = 0 (after square rooting both sides) w = 4

OpenStudy (anonymous):

@Hero great answer